OpenStudy (anonymous):

How much money should be deposited in an account that earns 9% compounded monthly so that it will accumulate to $10,000 in three years? Round to the nearest cent. :) THANK YOU FOR THE HELP!!

OpenStudy (anonymous):

@jim_thompson5910 Would you mind helping me with this one? :)

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) 10000 = x(1+0.09/12)^(12*3) solve for x

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

$7641.49? :)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sry i went afk for a bit

jimthompson5910 (jim_thompson5910):

you got it